four times the smaller is 57 more than the larger. the larger is 32 less than 3 times the smaller. Find the numbers.
Answer by checkley79(3341)

4X=Y+57
Y=3X-32
4X=(3X-32)+57
4X=3X+25
4X-3X=25
X=25 ANS. FOR THE SMALLER NUMBER.
Y=3*25-32
Y=75-32
Y=43 ANS. FOR THE LARGER NUMBER.
PROOF:
4*25=43+57
100=100